OpenStudy (anonymous):

HELP PLEASE!!! =DD Jose finished his math assignment in 1/8 hours. Then he completed his history assignment in 2/3 hours. How much more time did Jose spend on his history assignment? Write your answer as a fraction in simplest form.

OpenStudy (anonymous):

its simply 2/3-1/8..can you do the simplification now??

OpenStudy (anonymous):

2/3 - 1/8 = 16/24 - 3/24 (find the LCD in order to subtract it easily) = 13/24 hours